### v1

# PIR Over Wireless Channels: Achieving Privacy With Public Responses

Abstract
This paper addresses the problem of private information retrieval (PIR) over an additive white Gaussian noise (AWGN) channel, considering the channel is public. In such settings, each server can eavesdrop on the channel, potentially compromising the user’s privacy. Previous works suggested joint coding–PIR schemes, ignoring the fact that communication over a practical wireless channel is public. To address this gap, we present a novel joint wiretap–PIR coding scheme that leverages lattice codes to exploit the channel’s additive properties. This scheme integrates wiretap coding and private retrieval techniques into a unified framework.
Index Terms:
Private Information Retrieval, Wiretap, Lattice Codes, MISO.
I Introduction
In the classical PIR problem, there are $N$ identical, non-communicating servers, each storing the same $M$ messages, and a user who wishes to privately retrieve the $i$-th message without revealing the index $i$ to the servers. The primary objective is to minimize the communication overhead between the user and the servers while adhering to privacy constraints.
The PIR problem has recently attracted the attention of information theory society, motivated by the challenge of characterizing the fundamental limits of this problem. The classical PIR capacity, with noiseless and orthogonal channels, is given by ${C_{PIR}=(1-1/N)(1-(1/N)^{M})}$ in [1]. This key finding encourages the information theory community to explore further variations of this issue. Since then, many extensions and generalizations of the problem have been examined. For instance, robust PIR considers scenarios where some databases may fail to respond, and $T$-private PIR ensures privacy even when any $T$ out of the $N$ databases collude, as studied in [2]. Another variation is PIR from byzantine databases, which accounts for the possibility of up to $B$ databases providing erroneous responses, either intentionally or unintentionally [3].
While the PIR problem has been extensively studied over the past decades, most research has focused on simple communication channels characterized by orthogonal and noiseless links between the user and the database. The case of PIR over noisy orthogonal channels (NPIR) was examined in [4], which demonstrated that the channel coding required to mitigate channel errors is ”almost separable” from the retrieval scheme, depending only on the agreed traffic ratio. Additionally, [4] explored PIR over various types of MAC and showed that, unlike NPIR, the channel coding and retrieval scheme is not optimal in general. Building on this foundation, the authors in [5, 6, 7] investigated PIR over a block-fading Gaussian MAC.
In those papers, the authors did not refer to the fact that the channel model, which represents a practical wireless channel, is publicly available, hence making it possible for the servers to overhear the channel, which may cause a privacy leakage. This paper tackles the secrecy gap by utilizing methods developed for wiretap coding. The main idea is to prevent the servers from retaining any information related to the user from the observed signal.
In [8], the author extends the classical PIR problem to address scenarios involving colluding servers and eavesdroppers. This extended problem introduces additional privacy and secrecy constraints to the conventional PIR setting. Specifically, the user aims to retrieve a message without revealing the desired message index to any set of $T$ colluding, and an eavesdropper who can listen to the queries and answers of any $E$ servers but is prevented from learning any information about the messages. This problem is related to ours in the sense that it adds the secrecy constraint, and it can provide some insights. However, unlike [8], which assumes noiseless and orthogonal channels, our model assumes a MAC channel between the servers and the user, which presents a significant difference. The server could not collude in our model but can serve as an eavesdropper.
This paper is a direct continuation of [5], which tackles the problem of PIR over block–fading AWGN channel by considering that servers can also listen to the same channel used for their transmissions. We provide a joint wiretap-privacy-channel coding scheme to handle this. The resulting achievable rate is compared to the full channel capacity without any constraint when sending a single message. This comparison shows the redundancy of the suggested joint scheme, one has to tolerate to promise privacy.
II System Model Preliminaries
II-A Notational Conventions
Throughout the paper, we use boldface lowercase to refer to vectors, e.g., $\mathbf{h}\in\mathbb{R}^{L}$, and boldface uppercase to refer to matrices, e.g., $\mathbf{H}\in\mathbb{R}^{M\times L}$. For a vector $\mathbf{h}$, we write ${\|\mathbf{h}\|}$ for its Euclidean norm, i.e. ${\|\mathbf{h}\|}\triangleq\sqrt{\sum_{i}h_{i}^{2}}$. We denote by $\mathbf{e}_{i}$ the unit vector with $1$ at the $i$th entry and zero elsewhere. All logarithms in this paper are natural logarithms, and rates are in nats.
II-B System Model and Problem Statement
Consider the PIR problem over a public Gaussian wiretap channel with $N$ identical and non-communicating servers (transmitters), each equipped with one transmit antenna. Each server stores a set of messages $W_{1}^{M}=\{W_{1},W_{2},...,W_{M}\}$ of size $L$ each. The messages were drawn uniformly and independently from $\mathbb{F}_{p}^{L}$ where $p$ is assumed to be prime , i.e., in nats units, we have,
$H(W_{l})=L\log{p}\ \text{for}\ l=1,...,M,$ (1)
$H(W_{1}^{M})=ML\log{p}.$
As usual, the user wishes to privately retrieve the message $W_{i}$, where the index $i$ is assumed to be uniformly distributed on $[1,...,M]$, i.e., $i$ is a realization of $\theta\sim U[1,..M]$, while keeping $\theta$ secret from each server. However, unlike traditional PIR setups, herein, since the replies are over a wireless channel, it is important to keep $\theta$ private, also considering the public replies. Let $Q_{j}(\theta)$ denote a query that depends on the required random message index and is generated for the $j$th server. $Q_{j}(i)$ will be used to show the query’s dependence on the specific realization of the index, $i$. Accordingly, the user generates a set of $N$ queries $Q_{1}(i),Q_{2}(i),...,Q_{N}(i)$, one for each server, which are statistically independent with the messages (as those are not known to him). That is, we have
$I(W_{1}^{M};Q_{1}(\theta),...,Q_{N}(\theta))=0.$ (2)
We assume the servers receive the queries through independent control channels and cannot access each other’s queries or answers, a common assumption in the literature.
The $k$th server responds to its query with a message (or codeword) $\mathbf{x}_{k}(i)$ of size $n$, where $n$ is fixed for all responses. The reason for this is that we follow the usual Gaussian MAC setup in the literature [9], where a codeword is transmitted during $n$ channel uses. The $k$th server generates an answer based on the received query $Q_{k}(\theta)$ and the stored messages $W_{1}^{M}$. The answer, denoted by $\mathbf{x}_{k}(i)$, is a deterministic function of the messages and the query. We have
$H(\mathbf{x}_{k}(\theta)|W_{1}^{M},Q_{k}(\theta))=0.$ (3)
Because the server’s responses are transmitted over a public channel, each server can monitor other servers’ communications, potentially compromising user privacy. To ensure privacy, the server should not be able to infer any information about the desired message index given the queries, the entire database, and the received signal $\mathbf{w}$ at the curious server antenna. Thus, for each server $j$ the following privacy constraint must be satisfied,
$\begin{split}\text{[Secure - Privacy]}&\\ &\lim_{n\rightarrow\infty}I(\theta;Q_{j}(\theta),W_{1}^{M},\mathbf{w})=0,\hskip 5.69046pt\forall j\in\{1,...,N\}.\end{split}$ (4)
The databases are linked to the user via a block-memoryless fading AWGN channel (Figure 1). In this setup, the channel remains constant throughout the transmission of codewords of size $n$, and each block is independent of the others. Thus, over a transmission of $n$ symbols, the user and the curious server observe a noisy linear combination of the transmitted signals,
$\mathbf{y}=\sum_{k=1}^{N}h_{k}\mathbf{x}_{k}+\mathbf{n}_{y},\quad\mathbf{w}=\sum_{k=1}^{N}g_{k}\mathbf{x}_{k}+\mathbf{n}_{w}$ (5)
Here, $h_{k}\thicksim\mathcal{N}(0,1)$ and $g_{k}\thicksim\mathcal{N}(0,1)$ represent the real channel coefficients of the servers-user and servers-server channel, respectively, and $\mathbf{n}_{y}\thicksim\mathcal{N}(0,\sigma_{y}^{2}\mathbf{I}^{n\times n})$,$\mathbf{n}_{w}\thicksim\mathcal{N}(0,\sigma_{w}^{2}\mathbf{I}^{n\times n})$ are i.i.d. Gaussian noise observed by the user and the curious server, respectively.
Let us refine certain aspects of our model assumptions. We assume that explicit cooperation among servers is not permitted. Additionally, we assume a per-database power constraint, where all transmitting databases operate with a fixed power $P$, and power cannot be allocated differently to different servers. Thus, the transmitted codebook ${\cal C}$ must satisfy the average power constraint, i.e., $\operatorname{\mathbb{E}}\left[\|\mathbf{x}_{k}\|^{2}\right]\leq nP$.
We assume there is channel state information at the transmitter (CSIT) together with statistical information of the curious server, i.e., $\{h_{k}\}_{k=1}^{N}$,$\{g_{k}\}_{k=1}^{N}$ and $\sigma_{w}^{2}$ are known at the transmitter. Upon receiving the mixed response $\mathbf{y}$ from all the databases, the user decodes the required message $W_{i}$. Let $\widehat{W_{i}}$ denote the decoded message at the user and define the error probability for decoding a message as follows,
$P_{e}(L)\triangleq P_{r}(\widehat{W_{i}}\neq W_{i}).$ (6)
We require that $P_{e}(L)\rightarrow 0$ as $L$ tends to infinity.
Remark 1:
Note that the [Secure - Privacy] requirement, as defined above, is a special case that arises from the AWGN channel characteristic and is different from the usual conventions in the PIR filled. In our case, we will see that ensuring the Secure-Privacy constraint requires a strong secrecy notion from the wiretap channel.
Figure 1: System model of two servers connected to a user through an AWGN MAC channel while server $2$ is also aware of the transmission over the channel.
II-C Performance Metric
The PIR rate over MAC is defined as the total number of desired bits divided by the total number of channel uses [5, 7]. Specifically,
Definition 1:
Denote:
$R_{PIR}^{MAC}(n)\triangleq\frac{H(W_{i})}{n}=\frac{L\log{p}}{n},$
where $n$ represents the number of channel uses. The PIR capacity over AWGN MAC, denoted by $C_{PIR}^{MAC}$, is the supremum of $R_{PIR}^{MAC}(n)$, under which reliable communication is achievable, ensuring the privacy of the user’s will. That is, satisfying (2,3,4) and (6).
Neglecting the privacy constraint, our model matches with Multiple Input Single Output (MISO) channel with per-antenna power constraint model [10] hence this model will serve as an upper bound to our PIR capacity rate. In such a model, each transmit antenna has a separate power budget yet can fully cooperate with other antennas. The MISO sum capacity with per-antenna power constraint and fixed channel coefficients, which are globally known, is given by
$C_{SR}^{MISO}=\frac{1}{2}\log\left(1+P\left(\sum_{k=1}^{N}|h_{k}|\right)^{2}\right).$ (7)
We note that the MAC sum capacity is lower than $C_{SR}^{MISO}$ since, in the latter, the transmit antenna can fully cooperate with each other.
The following subsection introduces some preliminaries on the lattice, which we will need to describe and analyze our scheme.
II-D Lattice Code and Lattice Gaussian Distribution
It has been shown in [11, 12] that the full capacity of the point-to-point AWGN channel is achievable using lattice encoding and decoding. We now provide a brief background on lattice codes, which will be useful in the remainder of this paper.
An $n$-dimensional lattice $\Lambda$ is a discrete subgroup of the Euclidean space $\mathbb{R}^{n}$, which is closed under reflection and real addition. Recall the definition of an $n$-dimensional lattice [13].
Definition 2:
A non-degenerate $n$-dimensional lattice is defined by a set of $n$ linearly independent basis (column) vectors $\mathbf{g}_{1},...,\mathbf{g}_{n}$ in $\mathbb{R}^{n}$. The lattice is composed of all integral combinations of the basis vectors, i.e.,
$\Lambda=\left\{\lambda=\sum^{n}_{k=1}i_{k}\mathbf{g}_{k}\text{ : }i_{k}\in\mathbb{Z}\right\}=\left\{\lambda=G\cdot\mathbf{i}\text{ : }\mathbf{i}\in\mathbb{Z}^{n}\right\}$
where =$\mathbb{Z}=\{0,\pm 1,\pm 2,...\}$ is the set of integers, $\mathbf{i}=(i_{1},...,i_{n})^{t}$ is an $n$-dimensional integer (column) vector, and the $n\times n$ generator matrix $G$ is given by
$G=\left[\ \mathbf{g}_{1}\ |\ \mathbf{g}_{2}\ |\ ...\ |\ \mathbf{g}_{n}\ \right]$
The resulting lattice is denoted $\Lambda(G)$.
Definition 3 (Quantizer):
A lattice quantizer is a map, ${Q_{\Lambda}:\mathbb{R}^{n}\rightarrow\Lambda}$, that sends a point, $\mathbf{s}$, to the nearest lattice point in Euclidean distance. That is,
$Q_{\Lambda}(\mathbf{s})=\text{argmin}_{\lambda\in\Lambda}\left\|\mathbf{s-\lambda}\right\|.$ (8)
Definition 4 (Voronoi Region):
The fundamental Voronoi region, $\mathcal{V}$, of a lattice, $\Lambda$, is the set of all points in $\mathbb{R}^{n}$ that are closest to the zero vector compared to any other lattice point. That is, ${\mathcal{V}=\{\mathbf{s}:Q_{\Lambda}(\mathbf{s})=0\}}$.
Definition 5 (Modulus):
Let $[\mathbf{s}]\text{mod}\ \Lambda$ denote the quantization error of $\mathbf{s}\in\mathbb{R}^{n}$ with respect to the lattice $\Lambda$. That is,
$[\mathbf{s}]\text{mod}\ \Lambda=\mathbf{s}-Q_{\Lambda}(\mathbf{s}).$ (9)
For all $\mathbf{s,t}\in\mathbb{R}^{n}$ and $\Lambda_{c}\subseteq\Lambda_{f}$, the $\text{mod}\ \Lambda$ operation satisfies:
$[\mathbf{s}+\mathbf{t}]\text{mod}\ \Lambda=\big{[}[\mathbf{s}]\text{mod}\ \Lambda+\mathbf{t}\big{]}\text{mod}\ \Lambda$ (10)
$\left[Q_{\Lambda_{f}}(\mathbf{s})\right]\ \text{mod}\ \Lambda_{c}=\left[Q_{\Lambda_{f}}([\mathbf{s}]\ \text{mod}\ \Lambda_{c})\right]\ \text{mod}\ \Lambda_{c}$ (11)
$[a\mathbf{s}]\text{mod}\ \Lambda=[a[\mathbf{s}]\text{mod}\ \Lambda]\text{mod}\ \Lambda\quad\forall a\in\mathbb{Z}$ (12)
$\beta[\mathbf{s}]\text{mod}\ \Lambda=[\beta\mathbf{s}]\text{mod}\ \beta\Lambda\quad\forall\beta\in\mathbb{R}$ (13)
Next, we recall the definitions from [14] for Lattice Gaussian distributions, which will be needed to construct the PIR scheme. For $\sigma>0$ and $\mathbf{c}\in\mathbb{R}^{n}$, we define the Gaussian distribution of variance $\sigma^{2}$ centered at $\mathbf{c}\in\mathbb{R}^{n}$ as
$f_{\sigma,\mathbf{c}}(\mathbf{x})=\frac{1}{(\sqrt{2\pi}\sigma)^{n}}e^{-\frac{\|\mathbf{x}-\mathbf{c}\|^{2}}{2\sigma^{2}}}$
for all $\mathbf{x}\in\mathbb{R}^{n}$.
Definition 6 ($\Lambda$-Periodic Function):
$f_{\sigma,\Lambda}(\mathbf{x})=\sum_{\bm{\lambda}\in\Lambda}f_{\sigma,\bm{\lambda}}(\mathbf{x})=\frac{1}{(\sqrt{2\pi}\sigma)^{n}}\sum_{\bm{\lambda}\in\Lambda}e^{-\frac{\|\mathbf{x}-\bm{\lambda}\|^{2}}{2\sigma^{2}}}$
for all $\mathbf{x}\in\mathbb{R}^{n}$. Note that $f_{\sigma,\Lambda}$ restricted to the quotient $\mathbb{R}^{n}/\Lambda$ 111the set of all closets which in our case equivalent to the Voronoi region is a probability density.
Definition 7 (Discrete Gaussian):
discrete Gaussian distribution over $\Lambda$ centered at $\mathbf{c}\in\mathbb{R}^{n}$ and taking values over $\bm{\lambda}\in\Lambda$ defined as:
$D_{\Lambda,\sigma,\mathbf{c}}(\bm{\lambda})=\frac{f_{\sigma,\mathbf{c}}(\bm{\lambda})}{f_{\sigma,\Lambda}(\mathbf{c})},\quad\forall\bm{\lambda}\in\Lambda$
Lastly, we define the discrete Gaussian distribution over a coset of $\Lambda$, i.e., the shifted lattice $\Lambda-\mathbf{c}$ :
$D_{\Lambda-\mathbf{c},\sigma}(\bm{\lambda}-\mathbf{c})=\frac{f_{\sigma}(\bm{\lambda}-\mathbf{c})}{f_{\sigma,\Lambda}(\mathbf{c})}\quad\forall\bm{\lambda}\in\Lambda.$
III Main results
The following theorem presents an achievable strong secrecy PIR rate for the AWGN channel without fading, i.e., all the fading coefficients equal $1$. In our suggested scheme, we use the results provided in [14], which propose a wiretap lattice code that achieves strong secrecy and semantic security over the Gaussian wiretap channel (GWC). Their scheme achieves the secrecy capacity of the GWC up to a constant gap of $\frac{1}{2}$ nat. The first result is for $N=2$ servers.
Theorem 1:
For the $N=2$ databases public AWGN channel, the following strong secrecy PIR rate is achievable,
$R<\frac{1}{2}\log^{+}\left(\min\left\{\frac{\frac{1}{2}+SNR_{y}}{1+SNR_{w}},SNR_{y}\frac{\frac{1}{2}+SNR_{y}}{1+SNR_{y}}\right\}\right)-\frac{1}{2}$ (14)
where $SNR_{y}\frac{P}{\sigma^{2}_{y}}$, and $SNR_{w}=\frac{P}{\sigma^{2}_{w}}$.
The following Theorem extends the above result for $N>2$,
Theorem 2:
Consider the PIR problem with $N\geq 2$ databases over a public AWGN. Then, for any non-empty subsets of databases $\mathcal{S}_{1},\mathcal{S}_{2}$ satisfying ${\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset}$, $\mathcal{S}_{1}\cup\mathcal{S}_{2}\subseteq\{1,...,N\}$, the following PIR rate is achievable,
$R<\frac{1}{2}\log^{+}\left(\min\left\{\frac{\frac{1}{2}+SNR_{y}^{\prime}}{1+SNR_{w}^{\prime}},SNR_{y}^{\prime}\frac{\frac{1}{2}+SNR_{y}^{\prime}}{1+SNR_{y}^{\prime}}\right\}\right)-\frac{1}{2}$ (15)
where $SNR_{y}^{\prime}=\left\lfloor{\frac{N}{2}}\right\rfloor^{2}\frac{P}{\sigma^{2}_{y}}$, $SNR_{z}^{\prime}=\left\lfloor{\frac{N}{2}}\right\rfloor^{2}\frac{P}{\sigma^{2}_{w}}$.
Accordingly, the achievable rate scaling laws with the number of servers $N$ and the power $P$ are the same as those for the wiretap channel. The next theorem considers the PIR over public AWGN block–fading channel,
Theorem 3:
Consider the PIR problem with $N\geq 2$ databases over a public block-fading AWGN. Then, for any non-empty subsets of databases $\mathcal{S}_{1},\mathcal{S}_{2}$ satisfying ${\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset}$, $\mathcal{S}_{1}\cup\mathcal{S}_{2}\subseteq\{1,...,N\}$, the following PIR rate is achievable,
$R<\frac{1}{2}\log^{+}\left(\min_{i}\left\{\frac{\frac{1}{2}+SNR_{\tilde{y}}}{1+SNR_{\tilde{w}_{i}}},SNR_{\tilde{y}}\frac{\frac{1}{2}+SNR_{\tilde{y}}}{1+SNR_{\tilde{y}}}\right\}\right)-\frac{1}{2}$ (16)
where $\tilde{h}_{m}=\sum_{k\in\mathcal{S}_{m}}|h_{k}|$, $\tilde{g}_{m}=\sum_{k\in\mathcal{S}_{m}}g_{k}$, $SNR_{\tilde{y}}=\frac{\tilde{h}_{1}^{2}P}{\sigma_{y}^{2}}$, $SNR_{\tilde{w}_{1}}=\frac{\tilde{g}_{1}^{2}P}{\sigma_{w}^{2}}$, and $SNR_{\tilde{w}_{2}}=\frac{\left(\tilde{g}_{2}\frac{\tilde{h}_{1}}{\tilde{h}_{2}}\right)^{2}P}{\sigma_{w}^{2}}$.
IV A Joint PIR Scheme Over Public AWGN MISO channel
This section presents a PIR over the public AWGN MISO scheme.
IV-A A Joint Scheme for Secure PIR Over Non-Fading AWGN MAC With $N=2$ Servers
We start by analyzing the simple case with $N=2$ servers, as depicted in Figure 1. We assume that reciprocity between the channels. Therefore, it suffices to assume that only one server is eavesdropping on the channel. Let $\mathbf{y}$ and $\mathbf{z}$ denote the received inputs at the user and the malicious server (server number $2$), respectively. Over a transmission of $n$ symbols, the received input can be described as,
$\mathbf{y}(i)=\mathbf{x}_{1}+\mathbf{x}_{2}+\mathbf{n}_{y}$
and
$\mathbf{w}(i)=\mathbf{x}_{1}+\mathbf{x}_{2}+\mathbf{n}_{w}$
where $\mathbf{n}_{y}\sim\mathcal{N}(0,\sigma^{2}_{y}\mathbf{I}^{n\times n})$ and $\mathbf{n}_{w}\sim\mathcal{N}(0,\sigma_{w}^{2}\mathbf{I}^{n\times n})$ are the noises at the legitimate receiver and the eavesdropper, respectively. Assume, without loss of generality, that server $2$ can cancel its transmission. Thus, server $2$ obtains the following
$\mathbf{z}_{1}(i)=\mathbf{x}_{1}+\mathbf{n}_{w}.$
Note that the index $i$ in the received input denotes the private index of the desired message.
In the following, we provide an achievability scheme that proves Theorem 1.
IV-A1 Coding Scheme
Let $\Lambda_{w}\subset\Lambda_{f}$ be a nested chain of $n$-dimensional lattices, such that
$R=\frac{1}{n}\log|\Lambda_{f}/\Lambda_{w}|$
According to [15, Lemma 5], there exists a one-to-one mapping function, $\phi(\cdot)$, between a element $\mathbf{s}\in\mathbb{F}_{p}$, and a coset lattice point $\tilde{\bm{\lambda}}\in\Lambda_{f}/\Lambda_{w}$. For simplicity, we assume the coset representatives set is ${\Lambda_{f}\cap{\cal V}(\Lambda_{w})}$, namely:
$s\in\mathbb{F}_{p}\mapsto\bm{\lambda}=(\lambda_{1},...,\lambda_{n})\in\Lambda_{f}\cap{\cal V}(\Lambda_{w}).$
The set of coset representatives is not unique: One could choose $\bm{\lambda}\in\Lambda_{f}\cap\mathcal{R}(\Lambda_{w})$ for any fundamental region $\mathcal{R}(\Lambda_{w})$, not necessarily the Voronoi region $\mathcal{V}(\Lambda_{w})$.
Query: The user, which is interested in the message $W_{i}$, generates a random vector $\mathbf{b}$ of length $M$ such that each entry is either $1$ or 00, independently and with equal probability. Then, the user sends the following queries to the servers,
$Q_{1}(i)=\mathbf{b},\ \ Q_{2}(i)=-\mathbf{b}-\mathbf{e}_{i},\ \text{ if }b_{i}=0$ (17)
$Q_{1}(i)=\mathbf{b},\ \ Q_{2}(i)=-\mathbf{b}+\mathbf{e}_{i},\ \text{ if }b_{i}=1$
Answers: Upon receiving the queries, the servers construct their responses by computing linear combinations of the messages, where the combining coefficients are determined by the entries in $Q_{j}(i)$. That is,
$\mathbf{A}_{k}=\sum_{m=1}^{M}Q_{k}(i)W_{m},$ (18)
where $Q_{k}(m)$ is the $m$th entry of the vector $Q_{k}$. We note that $A_{k}=(a_{k}^{(1)},\dots,a_{k}^{(L)})\in\mathbb{F}_{p}^{L}$, and the scheme is focused on the transmission of a single symbol $a_{k}^{(m)}$, where $1\leq m\leq L$, from each answer. To construct the entire message, the servers must iterate this process across all $L$ symbols.
Without loss of generality, assume $m=1$, i.e., the servers wish to transmit the first symbol from each answer, that is, $a_{1}^{(1)}$ and $a_{2}^{(1)}$. Note that $\mathbf{A}_{1}+\mathbf{A}_{2}$ is equal to either $W_{i}$, or $-W_{i}$. This depends on the sign of $b_{i}$, which is known to the user. In the same way $a_{1}^{(m)}+a_{2}^{(m)}=\pm W_{i}^{(m)}$. To encode the symbols, each server maps the symbols to the codebook, namely, $\bm{\lambda}_{1}=\bm{\phi}(a_{1}^{(1)})$ and $\bm{\lambda}_{2}=\bm{\phi}(a_{2}^{(1)})$. Then, each server samples randomly and independently a lattice point from the relevant coset according to $D_{\Lambda_{w}+\bm{\lambda}_{i},\sqrt{P}}$ where $i\in\{1,2\}$ and $P$ is the transmission power. Equivalently, the servers transmissions are:
$\begin{split}&\mathbf{x_{1}}=\bm{\lambda}_{1}+\mathbf{r}_{1},\ \ \mathbf{r}_{1}\sim D_{\Lambda_{w},\sqrt{P},-\bm{\lambda}_{1}}\\ &\mathbf{x_{2}}=\bm{\lambda}_{2}+\mathbf{r}_{2},\ \ \mathbf{r}_{2}\sim D_{\Lambda_{w},\sqrt{P},-\bm{\lambda}_{2}}\end{split}$ (19)
Decode: To decode $\mathbf{v}\overset{\Delta}{=}\bm{\phi}(W_{i}^{(1)})$, the user computes the following,
$\hat{\mathbf{v}}=[Q_{\Lambda_{f}}(\alpha\mathbf{y})]\ \text{ mod }\Lambda_{w},$
where $\alpha=\frac{P}{P+\sigma^{2}_{y}}$ is the MMSE coefficient [14]. The above reduces to the Modulo-Lattice Additive Noise (MLAN) channel ([11]) as follows,
$\begin{split}\hat{\mathbf{v}}&=[Q_{\Lambda_{f}}(\alpha\mathbf{y})]\ \text{ mod}\Lambda_{w}\\ &=[Q_{\Lambda_{f}}(\alpha(\mathbf{x}_{1}+\mathbf{x}_{2}+\mathbf{n}_{y}))]\ \text{ mod }\Lambda_{w}\\ &=[Q_{\Lambda_{f}}(\mathbf{x}_{1}+\mathbf{x}_{2}+(\alpha-1)(\mathbf{x}_{1}+\mathbf{x}_{2})+\alpha\mathbf{n}_{y}))]\ \text{ mod }\Lambda_{w}\\ &=[Q_{\Lambda_{f}}(\mathbf{x}_{1}+\mathbf{x}_{2}+\mathbf{\tilde{n}}_{y}))]\ \text{ mod }\Lambda_{w}\end{split}$
where we denote the equivalent noise as ${\mathbf{\tilde{n}}_{y}\overset{\Delta}{=}(\alpha-1)(\mathbf{x}_{1}+\mathbf{x}_{2})+\alpha\mathbf{n}_{y}}$ where $\frac{1}{n}\operatorname{\mathbb{E}}[\left\|\mathbf{\tilde{n}}_{y}\right\|]=0$. Since $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ drawn independently from $D_{\Lambda_{w}+\bm{\lambda}_{i},\sqrt{P}}$, the second moment of ${\mathbf{\tilde{n}}_{y}}$ is given by ${\tilde{\sigma}_{y}^{2}=\frac{1}{n}\operatorname{\mathbb{E}}\left[\left\|\mathbf{n}_{eq}^{2}\right\|\right]=2(1-\alpha)^{2}P+\alpha^{2}\sigma_{y}^{2}}$ where we can optimize it on $\alpha$ getting $\alpha_{opt}=\frac{2P}{2P+\sigma_{y}^{2}}$ and the resulting optimal second moment is $\tilde{\sigma}^{2}_{y,opt}=\frac{2P\sigma_{y}^{2}}{2P+\sigma_{y}^{2}}$.
Using (11) we have,
$\begin{split}&[Q_{\Lambda_{f}}(\mathbf{x}_{1}+\mathbf{x}_{2}+\mathbf{\tilde{n}}_{y})]\ \text{ mod }\Lambda_{w}\\ &=[Q_{\Lambda_{f}}([\mathbf{x}_{1}+\mathbf{x}_{2}+\mathbf{\tilde{n}}_{y}]\text{ mod }\Lambda_{w})]\ \text{ mod }\Lambda_{w}\\ &=[Q_{\Lambda_{f}}([\mathbf{v}+\mathbf{r}_{1}+\mathbf{r}_{2}+\mathbf{\tilde{n}}_{y}]\text{ mod }\Lambda_{w})]\ \text{ mod }\Lambda_{w}\\ &\overset{(a)}{=}[Q_{\Lambda_{f}}([\mathbf{v}+\mathbf{\tilde{n}}_{y}]\text{ mod }\Lambda_{w})]\ \text{ mod }\Lambda_{w}\end{split}$
(a) since $\mathbf{r}_{1}\in\Lambda_{w}$ and $\mathbf{r}_{2}\in\Lambda_{w}$ together with (10). Therefore
$\hat{\mathbf{v}}=[Q_{\Lambda_{f}}(\mathbf{v}+\mathbf{\tilde{n}}_{y})]\ \text{ mod }\Lambda_{w}$
Then, from the AWGN-goodness of $\Lambda_{f}$ and the secrecy-goodness of $\Lambda_{w}^{(n)}$ (with respect to $\tilde{\sigma}_{w}=\frac{\sqrt{P}\sigma_{w}}{\sqrt{P^{2}+\sigma_{w}^{2}}}$), we have [11, 14]:
$\gamma_{\Lambda_{y}}\left(\tilde{\sigma}_{y}\right)=\frac{V\left(\Lambda_{y}\right)^{2/n}}{\tilde{\sigma}_{y}^{2}}>2\pi e,$ (20)
$\gamma_{\Lambda_{w}}\left(\tilde{\sigma}_{w}\right)=\frac{V\left(\Lambda_{w}\right)^{2/n}}{\tilde{\sigma}_{w}^{2}}<2\pi.$ (21)
We have that strong secrecy rates $R$ satisfying
$R=\frac{1}{n}\log\frac{V\left(\Lambda_{w}\right)}{V\left(\Lambda_{y}\right)}<\frac{1}{2}\log\left(\frac{\frac{1}{2}+SNR_{y}}{1+SNR_{w}}\right)-\frac{1}{2}$ (22)
where $SNR_{y}=\frac{P}{\sigma_{y}^{2}}$,and $SNR_{y}=\frac{P}{\sigma_{w}^{2}}$.
Two extra conditions are required for the average power constraint and for the equivalent noise to be asymptotically Gaussian, that is
$V\left(\Lambda_{w}\right)^{2/n}<2\pi\frac{P}{\pi-1},\ V\left(\Lambda_{w}\right)^{2/n}<2\pi\frac{P}{1+1/\rho_{b}}$
which together with (20) yields:
$\begin{split}&R<\frac{1}{2}\log\left(\frac{\frac{1}{2}+\operatorname{SNR}_{y}}{\frac{\pi}{\pi-1/e}}\right)-\frac{1}{2}\\ &R<\frac{1}{2}\log\left(SNR_{y}\frac{\frac{1}{2}+SNR_{y}}{1+SNR_{y}}\right)-\frac{1}{2}\end{split}$ (23)
Thus, combining (22)-(23) together, we obtain Theorem 1.
Remark 2:
Although the servers generate $r_{1}$ and $r_{2}$ independently, the user can decode the message. That is because both belong to $\Lambda_{w}$.
Next, we demonstrate that the secrecy-privacy requirement (4) for the $j$-th server is fulfilled.
Lemma 1:
$I(\theta;Q_{g}(\theta),W_{1}^{M})=0,\quad g\in\{1,2\}$ (24)
Proof.
This lemma follows immediately from the fact that $Q_{g}(\theta)$ is uniformly distributed over $\{0,1\}$ or $\{0,-1\}$ independently of $\theta$ and $W_{1}^{M}$. Additionally, $W_{1}^{M}$ independent with $\theta$ and $Q_{g}(\theta)$. ∎
Lemma 2:
$\lim_{n\rightarrow\infty}I(\lambda_{1},\lambda_{2};\mathbf{w})=0$ (25)
Proof.
$\begin{split}&I(\lambda_{1},\lambda_{2};\mathbf{w})\\ &=I(\lambda_{1};\mathbf{w})+I(\lambda_{2};\mathbf{w}\mid\lambda_{1})\\ &=H(\lambda_{1})-H(\lambda_{1}\mid\mathbf{w})+H(\lambda_{2}\mid\lambda_{1})-H(\lambda_{2}\mid\lambda_{1},\mathbf{w})\\ &\overset{(a)}{\leq}H(\lambda_{1})-H(\lambda_{1}\mid\mathbf{w},\lambda_{2})+H(\lambda_{2})-H(\lambda_{2}\mid\lambda_{1},\mathbf{w})\\ &\overset{}{=}H(\lambda_{1})-H(\lambda_{1}\mid\lambda_{1}+n_{z},\lambda_{2})\\ &\hskip 56.9055pt+H(\lambda_{2})-H(\lambda_{2}\mid\lambda_{2}+n_{z},\lambda_{1})\\ &\overset{(b)}{=}H(\lambda_{1})-H(\lambda_{1}\mid\lambda_{1}+n_{z})+H(\lambda_{2})-H(\lambda_{2}\mid\lambda_{2}+n_{z})\\ &\overset{}{=}I(\lambda_{1};\mathbf{z}_{1})+I(\lambda_{2};\mathbf{z}_{2})\overset{n\rightarrow\infty}{\longrightarrow}0\end{split}$
where (a) follows since conditioning reduces entropy and since $\lambda_{1}$ is independent with $\lambda_{2}$. (b) follows again since $\lambda_{1}$ independent with $\lambda_{2}$. The last step follows from the strong secrecy implied by our scheme.
Now, we ready to prove the Secure - Privacy (21) requirement is fulfilled,
$\begin{split}&I(\theta;Q_{j}(\theta),W_{1}^{M},\mathbf{w})\\ &=I(\theta;Q_{j}(\theta),W_{1}^{M})+I(\theta;\mathbf{w}\mid Q_{j}(\theta),W_{1}^{M})\\ &\overset{(a)}{=}H(\mathbf{w}\mid Q_{j}(\theta),W_{1}^{M})-H(\mathbf{w}\mid Q_{j}(\theta),W_{1}^{M},\theta)\end{split}$
(a) from Lemma 1. Assume, without loss of generality, that $j=2$. Thus,
$\begin{split}&H(\mathbf{w}\mid Q_{2}(\theta),W_{1}^{M})-H(\mathbf{w}\mid Q_{2}(\theta),W_{1}^{M},\theta)\\ &=H(\mathbf{w}\mid Q_{2}(\theta),W_{1}^{M})-H(\mathbf{w}\mid Q_{2}(\theta),W_{1}^{M},\theta,\lambda_{1},\lambda_{2})\\ &\overset{(b)}{\leq}H(\mathbf{w})-H(\mathbf{w}\mid\lambda_{1},\lambda_{2})\overset{n\rightarrow\infty}{\longrightarrow}0\end{split}$
(b) is due to the following Markov chain $\theta\leftrightarrow\left(Q_{1}(\theta),Q_{2}(\theta),W_{1}^{M}\right)\leftrightarrow\lambda_{1},\lambda_{2}\leftrightarrow\mathbf{w}$. The last step is due to Lemma 2. ∎
IV-B A Joint Scheme for Secure PIR Over Non-Fading AWGN MAC With $N>2$ Servers
This section extends the previously discussed results to scenarios involving more than $N=2$ servers. Specifically, over a transmission of $n$ symbols, the received signal at the user’s antenna is given by:
$\mathbf{y}(i)=\sum_{k=1}^{N}\mathbf{x}_{k}(i)+\mathbf{n}_{y},$ (26)
where $\mathbf{x}_{k}$ represents the signal transmitted by the $k$-th server, and $\mathbf{n}_{y}$ denotes the noise at the user. Similarly, the received signal at the malicious server’s antenna is described as:
$\mathbf{w}(i)=\sum_{k=1}^{N}\mathbf{x}_{k}(i)+\mathbf{n}_{w},$ (27)
where $\mathbf{n}_{w}$ represents the noise at the malicious server. It is assumed that the signals received by all other servers are subject to noisier channels. Consequently, ensuring security against the malicious server inherently provides protection against all other servers.
As proposed in [5], the user generates two queries (such as (17)) and divides the $N$ servers randomly into two groups of $\left\lfloor{\frac{N}{2}}\right\rfloor$ (in case $N$ is odd we can neglect one server). Each group subsequently constructs its answer based on the query assigned to it. The main distinction from the two-server scheme is the need to ensure that each group generates the same random lattice point (coset representative). To achieve coordination within each group, the user provides each group with a uniform binary vector. This vector is used to generate a sample from the relevant distribution [9] [^2] Generating a sample from a discrete distribution is a well-studied topic [9]. The expected number of fair bits required by an optimal algorithm to generate a random variable $X$ is less than $H(X)+2$., specifically $D_{\Lambda_{w}+\bm{\lambda}_{i},\sqrt{P}}$ where $i\in\{1,2\}$. Thus, the received input at the user and the malicious server is as follows:
$\begin{split}\mathbf{y}(i)=\left\lfloor{\frac{N}{2}}\right\rfloor\mathbf{x}_{1}(i)+\left\lfloor{\frac{N}{2}}\right\rfloor\mathbf{x}_{2}(i)+\mathbf{n}_{y}\\ \mathbf{w}(i)=\left\lfloor{\frac{N}{2}}\right\rfloor\mathbf{x}_{1}(i)+\left\lfloor{\frac{N}{2}}\right\rfloor\mathbf{x}_{2}(i)+\mathbf{n}_{w}\end{split}$ (28)
Without the loss of generality, assume that the malicious server belongs to group $2$ and can perfectly subtract the answers corresponding to its group (require knowledge about the number of servers participating in the transmission). Its objective becomes decoding $\mathbf{x}_{1}$ from the remaining received signals,
$\begin{split}\mathbf{z}_{1}(i)=\left\lfloor{\frac{N}{2}}\right\rfloor\mathbf{x}_{1}(i)+\mathbf{n}_{w}\end{split}$ (29)
This case is equivalent to the two servers scenario where the transmission power at each antenna is $(\left\lfloor{\frac{N}{2}}\right\rfloor)^{2}P$.
The only thing we have to verify is the decoding stage. Decode: To decode $\mathbf{v}\overset{\Delta}{=}\bm{\phi}(W_{i}^{(1)})$, the user computes the following,
$\hat{\mathbf{v}}=[Q_{\Lambda_{f}}(\alpha^{\prime}\mathbf{y})]\ \text{ mod }\Lambda_{w},$
where $\alpha^{\prime}=\alpha\frac{1}{\left\lfloor{\frac{N}{2}}\right\rfloor}$ and $\alpha$ will be given below. Thus,
$\begin{split}\hat{\mathbf{v}}&=[Q_{\Lambda_{f}}(\alpha^{\prime}\mathbf{y})]\ \text{ mod }\Lambda_{w}\\ &=[Q_{\Lambda_{f}}(\alpha(\mathbf{x}_{1}+\mathbf{x}_{2}+\frac{1}{\left\lfloor{\frac{N}{2}}\right\rfloor)}\mathbf{n}_{y})]\ \text{ mod }\Lambda_{w}\\ &=[Q_{\Lambda_{f}}(\mathbf{x}_{1}+\mathbf{x}_{2}+(\alpha-1)(\mathbf{x}_{1}+\mathbf{x}_{2})+\alpha\frac{1}{\left\lfloor{\frac{N}{2}}\right\rfloor)}\mathbf{n}_{y}))]\ \text{ mod }\Lambda_{w}\\ &[Q_{\Lambda_{f}}(\mathbf{x}_{1}+\mathbf{x}_{2}+\mathbf{\tilde{n}}_{y}))]\ \text{ mod }\Lambda_{w}\end{split}$
where we denote the equivalent noise as ${\mathbf{\tilde{n}}_{y}\overset{\Delta}{=}(\alpha-1)(\mathbf{x}_{1}+\mathbf{x}_{2})+\alpha^{\prime}\mathbf{n}_{y}}$ where $\frac{1}{n}\operatorname{\mathbb{E}}[\left\|\mathbf{\tilde{n}}_{y}\right\|]=0$. Since $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ drawn independently from $D_{\Lambda_{w}+\bm{\lambda}_{i},\sqrt{P}}$, the second moment of ${\mathbf{\tilde{n}}_{y}}$ is given by ${\tilde{\sigma}_{y}^{2}=\frac{1}{n}\operatorname{\mathbb{E}}\left[\left\|\mathbf{n}_{eq}^{2}\right\|\right]=2(1-\alpha)^{2}P+\alpha^{2}\left\lfloor{\frac{N}{2}}\right\rfloor^{-2}\sigma_{y}^{2}}$ where we can optimize it on $\alpha$ getting $\alpha_{opt}=\frac{2P}{2P+\left\lfloor{\frac{N}{2}}\right\rfloor^{-2}\sigma_{y}^{2}}$ and the resulting optimal second moment is $\tilde{\sigma}^{2}_{y,opt}=\frac{2P\left\lfloor{\frac{N}{2}}\right\rfloor^{-2}\sigma_{y}^{2}}{2P+\left\lfloor{\frac{N}{2}}\right\rfloor^{-2}\sigma_{y}^{2}}$.
Using (11) we have,
$\begin{split}&[Q_{\Lambda_{f}}(\mathbf{x}_{1}+\mathbf{x}_{2}+\mathbf{\tilde{n}}_{y})]\ \text{ mod }\Lambda_{w}\\ &=[Q_{\Lambda_{f}}([\mathbf{x}_{1}+\mathbf{x}_{2}+\mathbf{\tilde{n}}_{y}]\text{ mod }\Lambda_{w})]\ \text{ mod }\Lambda_{w}\\ &=[Q_{\Lambda_{f}}([\mathbf{v}+\mathbf{r}_{1}+\mathbf{r}_{2}+\mathbf{\tilde{n}}_{y}]\text{ mod }\Lambda_{w})]\ \text{ mod }\Lambda_{w}\\ &\overset{(a)}{=}[Q_{\Lambda_{f}}([\mathbf{v}+\mathbf{\tilde{n}}_{y}]\text{ mod }\Lambda_{w})]\ \text{ mod }\Lambda_{w}\end{split}$
(a) since $\mathbf{r}_{1}\in\Lambda_{w}$ and $\mathbf{r}_{2}\in\Lambda_{w}$ together with (10). Therefore
$\hat{\mathbf{v}}=[Q_{\Lambda_{f}}(\mathbf{v}+\mathbf{\tilde{n}}_{y})]\ \text{ mod }\Lambda_{w}$
Finally, following the same steps as in the two servers, as $n\rightarrow\infty$, we have that strong secrecy rate satisfies:
$R<\frac{1}{2}\log^{+}\left(\min\left\{\frac{\frac{1}{2}+SNR_{\tilde{y}}}{1+SNR_{\tilde{w}}},SNR_{y}^{\prime}\frac{\frac{1}{2}+SNR_{\tilde{y}}}{1+SNR_{\tilde{y}}}\right\}\right)-\frac{1}{2}$
where $SNR_{\tilde{y}}=\left\lfloor{\frac{N}{2}}\right\rfloor^{2}\frac{P}{\sigma^{2}_{y}}$ and $SNR_{\tilde{w}}=\left\lfloor{\frac{N}{2}}\right\rfloor^{2}\frac{P}{\sigma^{2}_{w}}$.
IV-C A Joint Scheme for PIR Over Public Block-Fading AWGN MAC with $N>2$ Servers
In this section, we consider a block-fading AWGN MAC. As discussed in the previous section, the transmissions here are affected not only by noise but also by attenuation due to the channel’s fading characteristics. The core concept of the proposed scheme involves dividing the servers into two groups and exploiting the channel characteristics to cancel the messages between them, leaving only the desired message at the receiver. The primary challenge is ensuring that the responses from both groups are properly aligned. To address this, the group with the higher channel gain reduces its transmission power to match the overall power of the two groups, facilitating alignment at the receiver.
IV-C1 Coding Scheme
Query: The user, which is interested in the message $W_{i}$, generates a random vector $\mathbf{b}$ of length $M$ such that each entry is either $1$ or 00, independently and with equal probability. Then, the user divides the databases into two non-intersecting subsets, denoted as $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, for which he sends the query $Q_{1}(i)$ to each member in $\mathcal{S}_{1}$ and $Q_{2}(i)$ to each member in $\mathcal{S}_{2}$. The queries are given as follows
$Q_{1}(i)=\mathbf{b},\ \ Q_{2}(i)=-\mathbf{b}-\mathbf{e}_{i},\ \text{ if }b_{i}=0$ (30)
$Q_{1}(i)=\mathbf{b},\ \ Q_{2}(i)=-\mathbf{b}+\mathbf{e}_{i},\ \text{ if }b_{i}=1$
Answers: Upon receiving the queries, the servers construct their responses by computing linear combinations of the messages, where the combining coefficients are determined by the entries in $Q_{j}(i)$. That is,
$\mathbf{A}_{k}=\sum_{m=1}^{M}Q_{k}(m)W_{m},$ (31)
where $Q_{k}(m)$ is the $m$th entry of the vector $Q_{k}$. We note that $A_{k}=(a_{k}^{(1)},\dots,a_{k}^{(L)})\in\mathbb{F}_{p}^{L}$, and the scheme is focused on the transmission of a single symbol $a_{k}^{(m)}$, where $1\leq m\leq L$, from each answer. To construct the entire message, the servers must iterate this process across all $L$ symbols.
Without any loss of generality, assume $m=1$, i.e., the servers wish to transmit the first symbol from each answer, that is $a_{1}^{(1)}$ and $a_{2}^{(1)}$. Note that $\mathbf{A}_{1}+\mathbf{A}_{2}$ is equal to either $W_{i}$, or $-W_{i}$. This depends on the sign of $b_{i}$, which is known to the user. In the same way $a_{1}^{(m)}+a_{2}^{(m)}=\pm W_{i}^{(m)}$. To encode the symbols, each server maps the symbols to the codebook, namely, $\bm{\lambda}_{1}=\bm{\phi}(a_{1}^{(1)})$ and $\bm{\lambda}_{2}=\bm{\phi}(a_{2}^{(1)})$. Then, each server samples randomly and independently a lattice point from the relevant coset according to $D_{\Lambda_{w}+\bm{\lambda}_{i},\sqrt{P}}$ where $i\in\{1,2\}$ and $P$ is the transmission power. Equivalently, the $i$-th server which belong to $\mathcal{S}_{g}$, $g\in\{1,2\}$ transmits:
$\begin{split}&\mathbf{x_{g}}=\bm{\lambda}_{g}+\mathbf{r}_{g},\ \ \mathbf{r}_{g}\sim D_{\Lambda_{w},\sqrt{P},-\bm{\lambda}_{g}}\end{split}$ (32)
The scheme requires that each group will generate the same sample. As in the non-fading scheme, the user is required to provide each group with a uniform binary vector. This vector will be used to generate a sample from the relevant distribution.
Over a transmission of $n$ symbols, the received signal at the user’s antenna and the malicious server (we consider the one with the best channel gain), respectively, is given by:
$\mathbf{y}(i)=\sum_{k\in\mathcal{S}_{1}}|h_{k}|\mathbf{x}_{1}(i)+\sum_{k\in\mathcal{S}_{2}}|h_{k}|\mathbf{x}_{2}(i)+\mathbf{n}_{y}$ (33)
$\mathbf{w}(i)=\sum_{k\in\mathcal{S}_{1}}|g_{k}|\mathbf{x}_{1}(i)+\sum_{k\in\mathcal{S}_{2}}|g_{k}|\mathbf{x}_{2}(i)+\mathbf{n}_{w},$ (34)
where $\mathbf{x}_{k}$ represents the signal transmitted by the $k$-th server, and $\mathbf{n}_{y}\sim\mathcal{N}(0,\sigma_{y}^{2}\mathbf{I}^{n\times n})$ and $\mathbf{n}_{w}\sim\mathcal{N}(0,\sigma_{w}^{2}\mathbf{I}^{n\times n})$ are an i.i.d., Gaussian noise, and $h_{k}\sim\mathcal{N}(0,1)$, $g_{k}\sim\mathcal{N}(0,1)$ are the real channel coefficients. Assuming full CSI, each server can adjust its transmission by multiplying its answer by $-1$ if necessary to ensure that the responses sum up according to the absolute values of the fading coefficients.
Assume, without loss of generality, $\tilde{h}_{1}\leq\tilde{h}_{2}$. Thus, in order to align the transmission, the servers that belong to $\mathcal{S}_{2}$ will reduce the transmission and transmit the following:
$\mathbf{x_{2}}^{\prime}=\frac{\tilde{h}_{1}}{\tilde{h}_{2}}\mathbf{x_{2}}$ (35)
where $\tilde{h}_{m}=\sum_{k\in\mathcal{S}_{m}}|h_{k}|$ and $\tilde{g}_{m}=\sum_{k\in\mathcal{S}_{m}}g_{k}$ (notice that only $h_{k}$ coefficients are with absolute value), which results in
$\begin{split}&\mathbf{y}(i)=\tilde{h}_{1}\mathbf{x}_{1}(i)+\tilde{h}_{2}\mathbf{x}^{\prime}_{2}(i)+\mathbf{n}_{y}=\tilde{h}_{1}(\mathbf{x}_{1}(i)+\mathbf{x}_{2}(i))+\mathbf{n}_{y}\\ &\mathbf{w}(i)=\tilde{g}_{1}\mathbf{x}_{1}(i)+\tilde{g}_{2}\mathbf{x}^{\prime}_{2}(i)+\mathbf{n}_{w}=\tilde{g}_{1}\mathbf{x}_{1}(i)+\tilde{g}_{2}\frac{\tilde{h}_{1}}{\tilde{h}_{2}}\mathbf{x}_{2}(i)+\mathbf{n}_{w}\end{split}$ (36)
Note that the user may add additional information to the query, informing the database which group the database belongs to and the factor to be multiplied before transmission.
Decode: To decode $\mathbf{v}\overset{\Delta}{=}\bm{\phi}(W_{i}^{(1)})$, the user computes the following,
$\hat{\mathbf{v}}=[Q_{\Lambda_{f}}(\alpha\frac{1}{\tilde{h}_{1}}\mathbf{y})]\ \text{ mod }\Lambda_{w},$
where $\alpha$ where $0\leq\alpha\leq 1$ will be determined later. The above reduces to the Modulo-Lattice Additive Noise (MLAN) channel ([11]) as follows,
$\begin{split}\hat{\mathbf{v}}&=[Q_{\Lambda_{f}}(\alpha\frac{1}{\tilde{h}_{1}}\mathbf{y})]\ \text{ mod }\Lambda_{w}\\ &=[Q_{\Lambda_{f}}(\alpha(\mathbf{x}_{1}+\mathbf{x}_{2}+\frac{1}{\tilde{h}_{1}}\mathbf{n}_{y}))]\ \text{ mod }\Lambda_{w}\\ &=[Q_{\Lambda_{f}}(\mathbf{x}_{1}+\mathbf{x}_{2}+(\alpha-1)(\mathbf{x}_{1}+\mathbf{x}_{2})+\alpha\frac{1}{\tilde{h}_{1}}\mathbf{n}_{y}))]\ \text{ mod }\Lambda_{w}\\ &=[Q_{\Lambda_{f}}(\mathbf{x}_{1}+\mathbf{x}_{2}+\mathbf{\tilde{n}}_{y}))]\ \text{ mod }\Lambda_{w}\end{split}$
where we denote the equivalent noise as ${\mathbf{\tilde{n}}_{y}\overset{\Delta}{=}(\alpha-1)(\mathbf{x}_{1}+\mathbf{x}_{2})+\alpha\frac{1}{\tilde{h}_{1}}\mathbf{n}_{y}}$ where $\frac{1}{n}\operatorname{\mathbb{E}}[\left\|\mathbf{\tilde{n}}_{y}\right\|]=0$. Since $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ drawn independently from $D_{\Lambda_{w}+\bm{\lambda}_{i},\sqrt{P}}$, the second moment of ${\mathbf{\tilde{n}}_{y}}$ is given by ${\tilde{\sigma}_{y}^{2}=\frac{1}{n}\operatorname{\mathbb{E}}\left[\left\|\mathbf{n}_{eq}^{2}\right\|\right]=2(1-\alpha)^{2}P+\alpha^{2}\frac{1}{\tilde{h}_{1}^{2}}\sigma_{y}^{2}}$ where we can optimize it on $\alpha$ getting $\alpha_{opt}=\frac{2P}{2P+\frac{1}{\tilde{h}_{1}^{2}}\sigma_{y}^{2}}$ and the resulting optimal second moment is $\tilde{\sigma}^{2}_{y,opt}=\frac{2P\frac{1}{\tilde{h}_{1}^{2}}\sigma_{y}^{2}}{2P+\frac{1}{\tilde{h}_{1}^{2}}\sigma_{y}^{2}}$.
Using (11) we have,
$\begin{split}&[Q_{\Lambda_{f}}(\mathbf{x}_{1}+\mathbf{x}_{2}+\mathbf{\tilde{n}}_{y})]\ \text{ mod }\Lambda_{w}\\ &=[Q_{\Lambda_{f}}([\mathbf{x}_{1}+\mathbf{x}_{2}+\mathbf{\tilde{n}}_{y}]\text{ mod }\Lambda_{w})]\ \text{ mod }\Lambda_{w}\\ &=[Q_{\Lambda_{f}}([\mathbf{v}+\mathbf{r}_{1}+\mathbf{r}_{2}+\mathbf{\tilde{n}}_{y}]\text{ mod }\Lambda_{w})]\ \text{ mod }\Lambda_{w}\\ &\overset{(a)}{=}[Q_{\Lambda_{f}}([\mathbf{v}+\mathbf{\tilde{n}}_{y}]\text{ mod }\Lambda_{w})]\ \text{ mod }\Lambda_{w}\end{split}$
(a) since $\mathbf{r}_{1}\in\Lambda_{w}$ and $\mathbf{r}_{2}\in\Lambda_{w}$ together with (10). Therefore
$\hat{\mathbf{v}}=[Q_{\Lambda_{f}}(\mathbf{v}+\mathbf{\tilde{n}}_{y})]\ \text{ mod }\Lambda_{w}$
For the analysis of the curious server, we assume the worst-case scenario: the server has full CSI for the legitimate channels and the inter-server channels ($h_{k}$ and $g_{k}$ for all $k$). Hence, the curious server can cancel part of the messages according to its group. Thus, the curious server observes one of the following, $\mathbf{z}_{1}(i)=\tilde{g}_{1}\mathbf{x}_{1}(i)+\mathbf{n}_{w}$ or $\mathbf{z}_{2}(i)=\tilde{g}_{2}\frac{\tilde{h}_{1}}{\tilde{h}_{2}}\mathbf{x}_{2}(i)+\mathbf{n}_{w}$. We will take care of both cases. For the first case we choose secrecy-good $\Lambda_{w}^{(n)}$ with respect to
$\tilde{\sigma}_{z_{1}}=\frac{\sqrt{P}\sigma_{w}\frac{1}{\tilde{g}_{1}}}{\sqrt{P^{2}+\sigma_{w}^{2}\frac{1}{\tilde{g}_{1}^{2}}}},$
and for the second, we choose $\Lambda_{w}^{(n)}$ with respect to
$\tilde{\sigma}_{z_{2}}=\frac{\sqrt{P}\sigma_{w}\frac{1}{\tilde{g}_{2}}\frac{\tilde{h}_{2}}{\tilde{h}_{1}}}{\sqrt{P^{2}+\sigma_{w}^{2}\left(\frac{1}{\tilde{g}_{2}}\frac{\tilde{h}_{2}}{\tilde{h}_{1}}\right)^{2}}}.$
Let us define $SNR_{\tilde{y}}=\frac{\tilde{h}_{1}^{2}P}{\sigma_{y}^{2}}$, $SNR_{\tilde{w}_{1}}=\frac{\tilde{g}_{1}^{2}P}{\sigma_{z}^{2}}$, and $SNR_{\tilde{w}_{2}}=\frac{\left(\tilde{g}_{2}\frac{\tilde{h}_{1}}{\tilde{h}_{2}}\right)^{2}P}{\sigma_{z}^{2}}$.
Thus, from the AWGN-goodness of $\Lambda_{f}$ and the secrecy-goodness of $\Lambda_{w}^{(n)}$ (with respect to $\tilde{\sigma}_{z_{1}}$ and $\tilde{\sigma}_{z_{2}}$), [11, 14], we obtain Theorem 3.
We note that the suggested scheme for the block-fading AWGN channel does not impair the Secure - Privacy (21) requirement, even if the channel coefficients ($\mathbf{h}$ and $\mathbf{g}$) are globally known. Specifically, we have
$\begin{split}&I(\theta;Q_{j}(\theta),W_{1}^{M},\mathbf{w},\mathbf{h},\mathbf{g})\\ &=I(\theta;\mathbf{h},\mathbf{g})+I(\theta;Q_{j}(\theta),W_{1}^{M},\mathbf{w}\mid\mathbf{h},\mathbf{g})\\ &=I(\theta;Q_{j}(\theta),W_{1}^{M}\mid\mathbf{h},\mathbf{g})+I(\theta;\mathbf{w}\mid Q_{j}(\theta),W_{1}^{M},\mathbf{h},\mathbf{g})\\ &\overset{(a)}{=}I(\theta;Q_{j}(\theta),W_{1}^{M})+I(\theta;\mathbf{w}\mid Q_{j}(\theta),W_{1}^{M},\mathbf{h},\mathbf{g})\\ &\overset{(b)}{=}H(\mathbf{w}\mid Q_{j}(\theta),W_{1}^{M},\mathbf{h},\mathbf{g})-H(\mathbf{w}\mid Q_{j}(\theta),W_{1}^{M},\theta,\mathbf{h},\mathbf{g})\\ &\overset{(c)}{\leq}H(\mathbf{w}\mid\mathbf{h},\mathbf{g})-H(\mathbf{w}\mid Q_{2}(\theta),W_{1}^{M},\theta,\mathbf{h},\mathbf{g})\\ &\overset{}{=}I(\mathbf{w};Q_{2}(\theta),W_{1}^{M},\theta\mid\mathbf{h},\mathbf{g})\\ &\overset{}{\leq}I(\mathbf{w};Q_{2}(\theta),W_{1}^{M},\theta,\lambda_{1},\lambda_{2}\mid\mathbf{h},\mathbf{g})\\ &\overset{(d)}{=}I(\mathbf{w};\lambda_{1},\lambda_{2}\mid\mathbf{h},\mathbf{g})\\ &=I(\mathbf{w};\lambda_{1}\mid\mathbf{h},\mathbf{g})+I(\mathbf{w};\lambda_{2}\mid\mathbf{h},\mathbf{g},\lambda_{1})\\ &=H(\lambda_{1}\mid\mathbf{h},\mathbf{g})-H(\lambda_{1}\mid\mathbf{h},\mathbf{g},\mathbf{w})\\ &\quad+H(\lambda_{2}\mid\mathbf{h},\mathbf{g},\lambda_{1})-H(\lambda_{2}\mid\mathbf{h},\mathbf{g},\lambda_{1},\mathbf{w})\\ &\leq H(\lambda_{1})-H(\lambda_{1}\mid\mathbf{h},\mathbf{g},\mathbf{w},\mathbf{x}_{2})\\ &\quad+H(\lambda_{2})-H(\lambda_{2}\mid\mathbf{h},\mathbf{g},\lambda_{1},\mathbf{w},\mathbf{x}_{1})\\ &\overset{(e)}{=}H(\lambda_{1})-H(\lambda_{1}\mid\mathbf{z}_{1})+H(\lambda_{2})-H(\lambda_{2}\mid\lambda_{1},\mathbf{z}_{2})\\ &=I(\lambda_{1};\mathbf{z}_{1})+I(\lambda_{2};\mathbf{z}_{2})\overset{n\rightarrow\infty}{\longrightarrow}0\end{split}$
(a) since ($\mathbf{h},\mathbf{g}$) is independent with ($\theta,Q_{j}(\theta),W_{1}^{M}$) and (b) is due to Lemma 1. (c) assume without loss of generality that $j=2$. (d) is due to the Markov chain $\theta\leftrightarrow\left(Q_{1}(\theta),Q_{2}(\theta),W_{1}^{M}\right)\leftrightarrow\left(\lambda_{1},\lambda_{2},\mathbf{h},\mathbf{g}\right)\leftrightarrow\mathbf{w}$. (e) is due to the Markov chains $\lambda_{1}\leftrightarrow\mathbf{x}_{1}\leftrightarrow\mathbf{z}_{1}\leftrightarrow(\mathbf{h},\mathbf{g},\mathbf{x}_{2},\mathbf{w})$, and $\lambda_{2}\leftrightarrow\mathbf{x}_{2}\leftrightarrow\mathbf{z}_{2}\leftrightarrow(\mathbf{h},\mathbf{g},\mathbf{x}_{1},\mathbf{w})$. The last step follows from the strong secrecy implied by our scheme.